A department of education reported that in 2007, 66% of students enrolled in college or a trade school within 12 months of granswer B
Answer #1

Part b

We have

Margin of error = Z* sqrt(p̂*(1 – p̂)/n)

We are given

Number of items of interest = x = 236

Sample size = n = 320

Sample proportion = p̂ = x/n = 236/320 = 0.7375

Confidence level = 95%

Critical Z value = 1.96

(by using z-table)

Margin of error = Z* sqrt(p̂*(1 – p̂)/n)

Margin of error = 1.96* sqrt(0.7375*(1 - 0.7375)/320)

Margin of error = 0.048209

Margin of error = 0.048

The margin of error is 0.048.